# Quantum State Preparation with Optimal T-Count

### v1

@@ Appendix B Improved analysis of Low-Kliuchnikov-Schaeffer state preparation
Let $n\geq 0$ be an integer and $U_{1},U_{2},\ldots,U_{2^{n}}$ be arbitrary single-qubit unitaries. Define $(n+1)$-qubit unitary $U=\mathsf{diag}(U_{1},\ldots,U_{2^{n}})$. Then $U$ can be implemented up to error $\varepsilon$ by a Clifford+${T}$ circuit using $O\left(\sqrt{2^{n}\log(1/\varepsilon)}+\log(1/\varepsilon)\right)$ ${T}$ gates and ancillary qubits.
@@
Any $n$-qubit state can be prepared up to error $\varepsilon$ by a Clifford+${T}$ circuit starting with the all-zeros state using
$O\left(\sqrt{2^{n}\log(1/\varepsilon)}+n\log(1/\varepsilon)\right)\text{ ${T}$gates and }O\left(\sqrt{2^{n}\log(1/\varepsilon)}+\log(1/\varepsilon)\right)\text{ ancillary qubits.}$. (98)
@@
The state synthesis algorithm in [31] uses the Grover-Rudolph algorithm [18]: The first qubit is rotated to have the correct marginal. Then conditioned on the first qubit, the second qubit is rotated to have the correct marginal. This process is iterated until the $n$th qubit is rotated conditioned on the first $n-1$ qubits.
@@
and the total number of ${T}$ gates used is

### v2

@@ Appendix B Improved analysis of Low-Kliuchnikov-Schaeffer state preparation
Let $n\geq 0$ be an integer and $U_{1},U_{2},\ldots,U_{2^{n}}$ be arbitrary single-qubit unitaries. Define $(n+1)$-qubit unitary $U=\mathsf{diag}(U_{1},\ldots,U_{2^{n}})$. Then $U$ can be implemented up to error $\varepsilon$ by a Clifford+T{T} circuit using $O\left(\sqrt{2^{n}\log(1/\varepsilon)}+\log(1/\varepsilon)\right)$ T{T} gates and ancillary qubits.
@@
Any $n$-qubit state can be prepared up to error $\varepsilon$ by a Clifford+T{T} circuit starting with the all-zeros state using
$O\left(\sqrt{2^{n}\log(1/\varepsilon)}+n\log(1/\varepsilon)\right)\text{ ${T}$ gates and }O\left(\sqrt{2^{n}\log(1/\varepsilon)}+\log(1/\varepsilon)\right)\text{ ancillary qubits.}$ (98)
@@
The state synthesis algorithm in [31] uses the Grover-Rudolph algorithm [18]: The first qubit is rotated to have the correct marginal. Then conditioned on the first qubit, the second qubit is rotated to have the correct marginal. This process is iterated until the $n$th qubit is rotated conditioned on the first n1n-1 qubits.
@@
and the total number of T{T} gates used is